OpenStudy (anonymous):

Describe the graph of y = 3/(4x-12) as compared to the graph of y = 1/x Sorry to post something that is so obviously a test question as just my second post, but I've been confused about this type of thing for a while. The only problem I have is that I don't know what the 4x does, but the rest I understand.

OpenStudy (anonymous):

I know that 3 is a vertical stretch of the equation by 3, and that -12 is a shift 12 to the right, but the 4x just baffles me to no end. I've checked my textbook, and have found questions like it, but no examples that help at all.

jimthompson5910 (jim_thompson5910):

y = 3/(4x-12) y = 3/(4(x-3)) y = (3/4)/(x-3) compare that last equation to y = a/(x - h) + k

OpenStudy (anonymous):

Thank you so much. Though I'm glad both of you answered, jim definitely was the best answerer here. I was thinking along those lines at some point and just totally lost my drift, thank you so much for getting me back on track.

jimthompson5910 (jim_thompson5910):

yw